OpenStudy (anonymous):

My integral calculus is a little rusty. There was an exercise in the first lecture series where we calculated the average power by integrating the COS waveform over one cycle. I'm sorry I don't have the problem in front of me. Can anyone show me how this is done? I'm a little confused. For example: how do you integrate (cos(2*pi*60t))^2 over one cycle? I fill in more detail when I get home tonight if needed.

OpenStudy (anonymous):

let y=(cos(2*pi*60*t))^2 cos^2t=(1+cos2t)/2 hence y=(1+cos2(2*pi*60t))/2 refer my solving thus we obtain y=t/2+sin[240*pi*t]/(480*pi)